OpenStudy (anonymous):

solve and check the equation. 15b - 6 = 8b - 3b + 14

OpenStudy (anonymous):

15b-6=5b+14 collect like terms 15b-5b=14+6 10b=20 divide both sides by 10 10b/10=20/10 b=2

OpenStudy (ash2326):

We have \[15b - 6 = 8b - 3b + 14\] Let's simplify the right hand side, combine the like terms \[15b-6=5b+14\] add 6 to both the sides \[15b=5b+14+6\] \[15b=5b+20\] subtract both sides by 5b \[15b-5b=20\] \[10b=20\] Divide both sides by 10 \[b=20/10\] we get \[b=2\] now to check this, put b=2 in the original equation and see if you get the same value both sides. If you get then we're right otherwise we're wrong. Will you check this?